OpenStudy (anonymous):

whats x? x+20+10x=20+x

OpenStudy (anonymous):

0

OpenStudy (anonymous):

11x + 20 = 20 + x -x -x 12x + 20 = 20 -20 -20 12x = 0 /12 /12 x = 0